Order: Lorazepam (Ativan) 0.5mg IV. Supply: vial labeled Lorazepam 1mg/ml. Dilute with 1 mL sterile water...

  1. Order: Lorazepam (Ativan) 0.5mg IV. Supply: vial labeled Lorazepam 1mg/ml. Dilute with 1 mL sterile water to yield 1mg/2mL. What is the quantity to administer?
  2. Order: Infuse 10% lipids 500 mL over 10 hours. What is the quantity to administer per hour?
  3. Order: 500 mL D5W I.V. at 50 mL/hr. How many hours will the I.V. run?
Answer #1

Ans) 1) Order: Lorazepam (Ativan) 0.5mg IV.

Supply: vial labeled Lorazepam 1mg/ml.

Dilute with 1 mL sterile water to yield 1mg/2mL.

1 mg = 2 ml

0.5 mg = x

x= 2×0.5 = 1 ml

2) Infuse 10% lipids 500 mL over 10 hours.

Formula: ml/hr = Volume to be administered/ time in hr

= 500/10 = 50 ml/hr.

3) Order: 500 mL D5W I.V. at 50 mL/hr.

Formula: ml/hr = Total volume to be administered/ time in hr

50 = 500/ time

Time = 500/50 = 10 hrs.